# MAGIC: A Multi-Hop and Graph-Based Benchmark for Inter-Context Conflicts in Retrieval-Augmented Generation

### v1

@@ Abstract
Jungyeon Lee, Kangmin Lee, Taeuk Kim* Hanyang University, Seoul, Republic of Korea {jungyune,kangmin7434,kimtaeuk}@hanyang.ac.kr
@@ 1 Introduction
Lastly, we conduct extensive analyses using MAGIC (A Multi-Hop And Graph-Based Benchmark for Inter-Context Conflicts), a novel dataset constructed through our framework. MAGIC features complex inter-context conflict patterns, including simultaneous and multi-hop cases rarely seen in existing benchmarks. Experiments on MAGIC yield key insights into how LLMs perceive knowledge conflicts: (1) most models remain imperfect at detecting conflicts, especially when multi-hop reasoning is required; and (2) even when contradictions are detected, models often fail to localize the exact point of conflict.
@@ 2 Related Work
Figure 2: Overview of the proposed KG-based framework for benchmarking inter-context knowledge conflict detection. It comprises three steps: (1) Subgraph Extraction, (2) Knowledge Conflict Generation, and (3) KG-to-Text Conversion, with details listed in Section 3.
@@ 3.1 Subgraph Extraction
As the first step, we distill parts of a large-scale KG to build knowledge segments that serve as targets for inducing knowledge conflicts. Theoretically, any KG can be utilized; in this work, we employ Wikidata5M Wang et al. (2021). Wikidata5M consists of approximately 20 million triplets, covering various domains and knowledge structures. [^4] For the diversity and robustness of MAGIC, we preprocess Wikidata5M as follows. Entities hard to be functionally defined, e.g., emoticons and special symbols (4,000 in total), are removed. In addition, general concepts and nodes with too many connections—e.g., ‘human’ and ‘United States’—are excluded. The 30 most connected nodes are filtered out.
@@ 3.2 Knowledge Conflict Generation
In the second phase, the goal is to perturb and modify extracted subgraphs to create counterparts that contradicts the original. To this end, we leverage LLMs with strong reasoning capabilities, expecting them to generate plausible and creative candidates that introduce knowledge conflicts within a given context. However, naïvely using such models does not guarantee success, as they are inherently imperfect at recognizing knowledge conflicts. We therefore propose a method to guide LLMs in reliably generating contradictory facts.
@@
Each conflict type is further divided into 1-conflict and N-conflict cases, based on the number of contradictions observed between two given contexts, yielding four distinct types illustrated in Figure 4. The numbers of data instances allocated to each category is shown in Table 1. This categorization enables fine-grained evaluation across varying reasoning depths and conflict complexities.
@@
Still, we observe that prompting without task demonstrations (i.e., real examples of knowledge conflicts) often falls short in generating diverse and logically complex conflicts, particularly in multi-hop scenarios. [^7] The model tends to repeat patterns, struggling to produce varied conflicts and often misaligning with common sense. As a solution, we adopt a few-shot prompting strategy, where the prompt includes three validated conflict examples per relation type. This encourages the model to move beyond simple entity or relation swaps. The final prompt template used in this process is shown in Figure 12, Appendix B.1.888While designed for multi-hop conflicts, the template can be readily adapted to single-hop settings. For N-conflict cases, the same graph is reused with different perturbations until the desired number of conflicts is achieved.
@@
Since LLMs are not inherently good at detecting knowledge conflicts, it may seem paradoxical to ask them to generate such cases. To alleviate this, we incorporate a human-in-the-loop process at two stages—before and after few-shot prompting—to improve dataset quality. First, human experts intervene during few-shot demonstration selection, manually filtering model-generated cases. Second, experts filter out trivial or incoherent outputs after conflict generation. [^9] Two researchers independently reviewed disjoint subsets using a shared guideline. See Appendix E for full details. Figure 5 depicts the entire workflow.
@@ 3.3 KG-to-Text Conversion
In addition, we validate data integrity by sampling and manually inspecting a subset of generated instances. Our manual inspection confirms that the conversion process achieves consistently high quality; further details are provided in Appendix B.2.
@@ 4 Experimental Setups
Prior work Jiayang et al. (2024); Hou et al. (2024) typically frames the task as a binary classification problem, relying on minimal prompts (see Appendix B.3). In contrast, we introduce a stepwise prompting strategy (Figure 16) designed to more fully probe the capabilities of LLMs for inter-context conflict detection.
  1. (1)
    Identification: LLMs determine whether a conflict exists between the given passages.
  2. (2)
    Explanation: If so, they specify how many conflicts are present and explain why, encouraging logical reasoning beyond surface cues.
  3. (3)
    Localization: LLMs pinpoint the exact sentences where conflicts occur, assessing their ability to locate contradiction sources.
@@
To account for the stochasticity of LLMs, all models perform three independent inference runs. We use two metrics for fine-grained evaluation, with scores averaged across all instances in each dataset. These metrics are manually computed by participating researchers, as automatic methods—such as LLM-as-a-judge—remain insufficiently reliable for this task. [^11] If they were, further investigation into knowledge conflict detection would be unnecessary.
@@
Table 2: ID scores (%) on three KC detection datasets. Lower scores indicate higher task complexity.
@@ 5 Experimental Results
LLMs tested on MAGIC consistently show lower ID and LOC scores compared to those on ECON and WikiContradict, with average scores decreasing by up to 15% and 19%, respectively. This indicates that models struggle more to identify conflicts in our dataset, and even when they do, they have difficulty pinpointing the exact portions where the conflict occurs.
@@
Table 2 shows that GPT-4o-mini achieves the highest accuracy and generalizes well to MAGIC. Mixtral performs the worst overall, while Haiku drops notably on MAGIC, suggesting difficulty with multi-hop reasoning. Llama shows a unique trend on MAGIC: it fails to respond in 47.4% of runs but achieves high ID scores when it does. Finally, o1 exhibits moderate performance, implying that conflict detection may depend on more than just LLMs’ reasoning ability.
@@
Table 3: LOC scores (%) on three KC detection datasets. Lower scores indicate higher complexity. We observe that models struggle more with pinpointing the exact location of a conflict than with detecting its presence.
@@
Table 3 shows trends similar to Table 2, with GPT-4o-mini consistently achieving the best LOC scores across all datasets. Qualitative analysis reveals that o1 often takes a conservative stance, frequently predicting no conflict in ambiguous cases,121212e.g., (A consists of B) vs. (A has C), (B not part of C). which leads to missed conflicts—particularly those requiring multi-hop thinking (see Table 8). [^13] This outcome is unexpected, given o1’s strong reasoning abilities in math and coding. One possible explanation is the exclusion of explicit reasoning instructions such as “think step by step”. We leave further investigation on this as future work. Mixtral’s low LOC score largely stems from poor initial conflict identification, reducing the number of evaluated localization cases. Llama shows moderate performance but produces overly long outputs (1121.5 tokens vs. 631.1 for others), often including irrelevant content. This suggests over-reliance on instructions, leading to over-detection and reduced localization precision.
Figure 7: Average performance of LLMs on the 5 most predictive and 5 least predictive relations. This analysis focuses soly on the single-conflict subset of MAGIC.
@@
Meanwhile, a higher number of conflicts reflects a stronger contradiction between the two contexts, making conflict detection easier for models. [^14] A similar trend is also observed in ECON Jiayang et al. (2024). See Appendix G for details. In other words, while more conflicts facilitate identification, they also complicate precise localization. When multiple conflicts occur, pinpointing all specific conflicting sentences becomes more difficult, leading to lower LOC scores.
@@
As MAGIC spans diverse topics and relations in KGs, it enables analysis at both the domain and relation levels. [^15] Detailed domain-level statistics are reported in Table 9. In the domain-specific evaluation, we find that most LLMs perform relatively well on the Class/Concept domain, likely because it contains clear and well-defined relations (e.g., subclass of, different from) that are readily recognizable. [^16] Still, exceptions exist depending on the used model. In contrast, performance on the Organization domain varies widely—from around 75% (GPT-4o-mini, Llama) to below 15% (Haiku, Mixtral)—indicating that models differs in handling hierarchical relations.
Meanwhile Figure 7 highlights the top five most and least predictive relations across all LLMs. Conflicts involving captain and mother are relatively easy for models, while work location and father pose greater challenges. More representative examples are provided in Table 13 in the Appendix.
@@ 6 Discussion
Comparison of challenging subsets
We conduct a comparative study between the challenging subset of WikiContradict (Implicit) and the corresponding subset from our dataset (Multi-Hop in MAGIC). [^17] WikiContradict consists of explicit and implicit conflicts; the latter are considered more challenging due to their subtlety. Figure 9 reports that MAGIC proves more difficult for LLMs than WikiContradict. MAGIC yields an ID score up to 15% lower and a LOC score up to 23% lower compared to WikiContradict. This highlights MAGIC’s intrinsic complexity, setting it apart from existing datasets.
@@
Figure 11: ID scores of two LLMs on known (O) vs. unknown (X) instances, based on their parametric knowledge. Models demonstrate stronger conflict detection when the relevant knowledge is already embedded.
@@
To address this, we explore an underexamined setup by splitting a subset of MAGIC—instances with 1- and 2-conflicts—into two groups based on the estimated presence of parametric knowledge in LLMs. Concretely, we approximate the existence of parametric knowledge for a given triplet by posing a converted verification question—for example, ‘Is Barack Obama Sr. the father of Barack Obama? We label a triplet as known if the model provides the correct answer in at least 4 out of 5 attempts, and as unknown if it succeeds in no more than 1.
As shown in Figure 11, both GPT-4o-mini and Llama 3.1 achieve higher ID scores on known instances. These findings suggest that models may be more effective at detecting conflicts when they already possess the relevant factual knowledge. While this approach does not provide a rigorous measure of internal knowledge, it offers a coarse-grained perspective. We remain a more systematic investigation for future work.
@@ 7 Conclusion
While MAGIC offers a novel benchmark for evaluating knowledge conflict detection, particularly inter-context conflict, it still presents several areas for improvement. First, MAGIC is constructed using Wikidata-based knowledge graphs. Incorporating additional sources—such as DBpedia, YAGO, or domain-specific knowledge graphs—could enhance its robustness and broaden its applicability. A promising direction is to align semantically equivalent relations across these graphs to ensure consistency and coverage. In addition, localization evaluation currently relies on human judgment in this work, developing more automated approaches could enable more fine-grained evaluation in future work. Addressing these limitations in future work will help enhance the robustness, scalability, and applicability of knowledge conflict detection in large-scale AI systems.
@@ Appendix A Details of Selected Relation Lists per Domain
Appendix B Prompts used in MAGIC
B.1 Prompt for Knowledge Conflict Generation
@@
Figure 12: Prompt for generating multi-hop conflicts.
B.2 Prompt for KG-to-Text Conversion
@@
Figure 13: Prompt for KG-to-Text Conversion.
Figure 14: Prompt for KG-to-Text Verification.
B.3 Prompt for Knowledge Conflict Detection
@@
Appendix C Linguistic quality of MAGIC
Since the contexts in MAGIC are generated by LLMs, assessing their linguistic quality is important to ensure their naturalness and realism. To this end, we conducted an evaluation using Claude 3.7 Sonnet. For a fair comparison, we sampled 155 contexts from each dataset (WikiContradict, ECON, MAGIC) to match overall length distributions. Each context was rated on a 0~5 scale, and the evaluation prompt is provided in Figure 17 and 18.
As shown in Table 6, MAGIC is only 0.32 lower in naturalness than WikiContradict, despite the latter being based on real Wikipedia text. It also surpasses ECON in realism, indicating our dataset generates plausible, real-world-like scenarios. These results demonstrate that MAGIC serves as a reliable benchmark for evaluating knowledge conflict detection.
@@ Appendix C Linguistic quality of MAGIC
[…]
Table 6: Naturalness and realism scores of KC datasets.

### v3

@@ Abstract
Jungyeon Lee1 †, Kangmin Lee1 2 †, Taeuk Kim1 * 1Hanyang University, 2KT Corporation {jungyune,kimtaeuk}@hanyang.ac.kr, lee.kangmin@kt.com
@@ 1 Introduction
Lastly, we conduct extensive analyses using MAGIC (A Multi-Hop And Graph-Based Benchmark for Inter-Context Conflicts), a novel dataset constructed through our framework. [^2] Available at: https://github.com/HYU-NLP/MAGIC. MAGIC features complex inter-context conflict patterns, including simultaneous and multi-hop cases rarely seen in existing benchmarks. Experiments on MAGIC yield key insights into how LLMs perceive knowledge conflicts: (1) most models remain imperfect at detecting conflicts, especially when multi-hop reasoning is required; and (2) even when contradictions are detected, models often fail to localize the exact point of conflict.
@@ 2 Related Work
Figure 2: Overview of the proposed KG-based framework for benchmarking inter-context knowledge conflict detection. It comprises three steps: (1) Subgraph Extraction, (2) Knowledge Conflict Generation, and (3) KG-to-Text Conversion, with details listed in §3.
@@ 3.1 Subgraph Extraction
As the first step, we distill parts of a large-scale KG to build knowledge segments that serve as targets for inducing knowledge conflicts. Theoretically, any KG can be utilized; in this work, we employ Wikidata5M Wang et al. (2021). Wikidata5M consists of approximately 20 million triplets, covering various domains and knowledge structures. [^5] For the diversity and robustness of MAGIC, we preprocess Wikidata5M as follows. Entities hard to functionally define, e.g., emoticons and special symbols (4,000 in total), are removed. In addition, general concepts and nodes with too many connections—e.g., ‘human’ and ‘United States’—are excluded. The 30 most connected nodes are filtered out.
@@ 3.2 Knowledge Conflict Generation
In the second phase, the goal is to perturb and modify extracted subgraphs to create counterparts that contradict the original. To this end, we leverage LLMs with strong reasoning capabilities, expecting them to generate plausible and creative candidates that introduce knowledge conflicts within a given context. [^7] Beyond rule-based substitutions at the entity or relation level as commonly explored in prior work, we aim to induce more natural and contextually coherent conflicts using LLMs. However, naïvely using such models does not guarantee success, as they are inherently imperfect at recognizing knowledge conflicts. We therefore propose a method to guide LLMs in reliably generating contradictory facts.
@@
Each conflict type is further divided into 1-conflict and N-conflict cases, based on the number of contradictions observed between two given contexts, yielding four distinct types illustrated in Figure 4. The number of data instances allocated to each category is shown in Table 1. This categorization enables fine-grained evaluation across varying reasoning depths and conflict complexities.
@@
Still, we observe that prompting without task demonstrations (i.e., real examples of knowledge conflicts) often falls short in generating diverse and logically complex conflicts, particularly in multi-hop scenarios. [^9] The model tends to repeat patterns, struggling to produce varied conflicts and often misaligning with common sense. As a solution, we adopt a few-shot prompting strategy, where the prompt includes three validated conflict examples per relation type. This encourages the model to move beyond simple entity or relation swaps. The final prompt template used in this process is shown in Figure 5.101010While designed for multi-hop conflicts, the template can be readily adapted to single-hop settings. For N-conflict cases, the same graph is reused with different perturbations until the desired number of conflicts is achieved.
Figure 5: Prompt for generating multi-hop conflicts.
@@
Since LLMs are not inherently good at detecting knowledge conflicts, it may seem paradoxical to ask them to generate such cases. To alleviate this, we incorporate a human-in-the-loop process at two stages—before and after few-shot prompting—to improve dataset quality. First, human experts intervene during few-shot demonstration selection, manually filtering model-generated cases. Second, experts filter out trivial or incoherent outputs after conflict generation. [^11] Two researchers independently reviewed disjoint subsets using a shared guideline. See Appendix G for full details. Figure 6 depicts the entire workflow.
@@ 3.3 KG-to-Text Conversion
In addition, we validate data integrity by sampling and manually inspecting a subset of generated instances. Our manual inspection confirms that the conversion process achieves consistently high quality; further details are provided in Appendix B.
@@ 4 Experimental Setups
Prior work Jiayang et al. (2024); Hou et al. (2024) typically frames the task as a binary classification problem, relying on minimal prompts (see Appendix B). In contrast, we introduce a stepwise prompting strategy (refer to Figure 17 for details) designed to more fully probe the capabilities of LLMs for inter-context conflict detection.
@@
To account for the stochasticity of LLMs, all models perform three independent inference runs. We use two metrics for fine-grained evaluation, with scores averaged across all instances in each dataset. These metrics are manually computed by participating researchers, as automatic methods—such as LLM-as-a-judge—remain insufficiently reliable for this task. [^13] If they were, further investigation into knowledge conflict detection would be unnecessary. See Appendix H for details.
@@
Table 2: ID scores (%) on three KC detection datasets. Lower scores indicate higher task complexity. The lowest value in each row is shown in bold.
@@ 5 Experimental Results
LLMs tested on MAGIC consistently show lower ID and LOC scores compared to those on ECON and WikiContradict, with average scores decreasing by up to 10% and 17%, respectively. This indicates that models struggle more to identify conflicts in our dataset, and even when they do, they have difficulty pinpointing the exact portions where the conflict occurs.
@@
Table 2 shows that GPT-4o-mini achieves the highest accuracy and generalizes well to MAGIC. Mixtral performs the worst overall, while Haiku still lags behind larger models, suggesting difficulty with multi-hop reasoning. Llama shows a unique trend on MAGIC: it fails to respond in 47.4% of runs but achieves high ID scores when it does. Finally, o1 exhibits moderate performance, implying that conflict detection may depend on more than just LLMs’ reasoning ability.
@@
Table 3: LOC scores (%) on three KC detection datasets. Lower scores indicate higher complexity. We observe that models struggle more with pinpointing the exact location of a conflict than with detecting its presence. The lowest value in each row is shown in bold.
@@
Table 3 shows trends similar to Table 2, with GPT-4o-mini consistently achieving the best LOC scores across all datasets. Qualitative analysis reveals that o1 often takes a conservative stance, frequently predicting no conflict in ambiguous cases,141414e.g., (A consists of B) vs. (A has C), (B not part of C). which leads to missed conflicts—particularly those requiring multi-hop thinking (see Table 10). [^15] This outcome is unexpected, given o1’s strong reasoning abilities in math and coding. One possible explanation is the exclusion of explicit reasoning instructions such as “think step by step”. We leave further investigation on this as future work. Mixtral’s low LOC score largely stems from poor initial conflict identification, reducing the number of evaluated localization cases. Llama shows moderate performance but produces overly long outputs (1121.5 tokens vs. 631.1 for others), often including irrelevant content. This suggests over-reliance on instructions, leading to over-detection and reduced localization precision.
@@
Meanwhile, a higher number of conflicts reflects a stronger contradiction between the two contexts, making conflict detection relatively easier for models. In other words, while more conflicts facilitate identification, they also complicate precise localization. When multiple conflicts occur, pinpointing all the individual conflicting sentences becomes more difficult, thereby leading to lower LOC scores. A similar trend has also been observed in ECON in Appendix E supports this finding.
Figure 8: Average performance of LLMs on the 5 most predictive and 5 least predictive relations. This analysis focuses solely on the single-conflict subset of MAGIC.
@@
As MAGIC spans diverse topics and relations in KGs, it enables analysis at both the domain and relation levels. Detailed domain-level statistics are reported in Table 11. In the domain-specific evaluation, we find that most LLMs perform relatively well on the Class/Concept domain, likely because it contains clear and well-defined relations (e.g., subclass of, different from) that are readily recognizable. [^16] Still, exceptions exist depending on the used model. In contrast, performance on the Organization domain varies widely—from around 75% (GPT-4o-mini, Llama) to below 45% (Mixtral, Haiku, o1)—indicating that models differs in handling hierarchical relations.
Meanwhile, Figure 8 highlights the top five most and least predictive relations across all LLMs. Conflicts involving captain and mother are relatively easy for models, while work location and father pose greater challenges. More representative examples are provided in Table 17 in the Appendix.
@@ 6 Further Analysis and Discussion
Figure 10: Comparison of ID and LOC scores on 1-Single-Hop conflicts across three datasets.
Comparison with existing KC datasets
In addition to the results discussed in §5, we compare MAGIC with existing datasets by focusing on the 1-Single-Hop type—the only category shared across all datasets—to ensure a fair and consistent comparison. In ECON and WikiContradict, conflict categories are manually annotated. Figure 10 shows that even on relatively simple 1-Single-Hop conflicts, models perform worse on MAGIC than on ECON and WikiContradict, suggesting that our dataset poses challenges beyond surface-level contradiction.
We further conduct a comparative study between the challenging subset of WikiContradict (Implicit) and the corresponding subset from our dataset (Multi-Hop in MAGIC). [^17] WikiContradict consists of explicit and implicit conflicts; the latter are considered more challenging due to their subtlety. Figure 11 reports that MAGIC proves more difficult for LLMs than WikiContradict. MAGIC yields an ID score up to 9% lower and a LOC score up to 16% lower compared to WikiContradict. This highlights MAGIC’s intrinsic complexity, setting it apart from existing datasets.
@@
Figure 13: ID scores of two LLMs on known (O) vs. unknown (X) instances, based on their parametric knowledge. Models show slightly stronger conflict detection when the relevant knowledge is already embedded.
@@
To address this, we explore an underexamined setup by splitting a subset of MAGIC—instances with 1- and 2-conflicts—into two groups based on the estimated presence of parametric knowledge in LLMs. Concretely, we approximate the existence of parametric knowledge for a given triplet by posing a converted verification question—for example, ‘Is Barack Obama Sr. the father of Barack Obama?’. We label a triplet as known if the model provides the correct answer in at least 4 out of 5 attempts, and as unknown if it succeeds in no more than 1.
As shown in Figure 13, both GPT-4o-mini and Llama 3.1 achieve slightly higher ID scores on instances with prior knowledge. These findings suggest that models may detect conflicts more effectively when they already possess the relevant factual knowledge, although a more systematic investigation is left for future work.
Impact of context order
To probe whether context order influences model behavior, we evaluate both the original MAGIC and a variant in which the two contexts in each instance are reversed.
As reported in Table 4, the ID scores remain largely consistent across models, with Mixtral 8x7B showing a slight increase and others nearly unchanged. These results indicate that context order does not significantly affect model performance on MAGIC, suggesting that our findings are robust to input ordering.
[…]
Table 4: Effect of context order on ID scores in MAGIC.
Linguistic quality of MAGIC
An essential requirement for MAGIC is that its conflict data be both realistic and stylistically similar to human-written text. To validate this, we use an LLM judge (Claude 3.7 Sonnet) to evaluate two aspects: (i) naturalness—whether the text is fluent and human-like—and (ii) realism—whether the scenario appears plausible in real-world contexts. We randomly sample 155 contexts from three datasets to match overall length distributions. Each context is rated on a 0–5 scale, with the detailed evaluation prompts provided in Figures 18 and 19.
As shown in Table 5, MAGIC achieves naturalness scores close to WikiContradict (-0.31), despite WikiContradict relying on real Wikipedia text. It further surpasses ECON in realism, showing that MAGIC generates scenarios that are both plausible and human-like. These results underscore MAGIC’s reliability as a benchmark for advancing research in knowledge conflict detection.
[…]
Table 5: Naturalness and realism scores for three KC datasets, evaluated by Claude 3.7 Sonnet.
@@ 7 Conclusion
While MAGIC provides a novel benchmark for evaluating knowledge conflict detection, particularly inter-context conflict, it still leaves room for improvement. First, MAGIC is built on Wikidata-based knowledge graphs; incorporating additional or domain-specific sources could improve robustness and applicability, and we provide preliminary examples from DBpedia and YAGO in the appendix. Future work also includes aligning semantically equivalent relations across graphs to ensure consistency and coverage. In addition, since localization evaluation currently relies on human judgment, developing automated approaches could enable a more fine-grained assessment.
Ethics Statement
We release MAGIC, a benchmark dataset for inter-context knowledge conflict detection. The dataset is constructed from Wikidata5M, a publicly available resource under permissive licenses, and contains no personally identifiable information. While we carefully filtered and validated the generated contexts, we cannot fully guarantee the absence of biased or potentially offensive content. MAGIC is released solely for scientific research, with the aim of fostering the development of systems that can reliably handle conflicting information and support safe deployment in real-world applications.
Acknowledgments
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-00558151). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2020-II201373, Artificial Intelligence Graduate School Program(Hanyang University)). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) under the artificial intelligence semiconductor support program to nurture the best talents (IITP-(2025)-RS-2023-00253914) grant funded by the Korea government(MSIT).
@@ Appendix A Details of Selected Relation Lists per Domain
Appendix B Detailed Prompts used in MAGIC
Prompt for knowledge conflict generation
@@ Appendix B Detailed Prompts used in MAGIC
Prompt for KG-to-Text conversion
@@
Figure 14: Prompt for KG-to-Text Conversion.
Figure 15: Prompt for KG-to-Text verification.
@@
Prompt for knowledge conflict detection
@@
Prompt for evaluating linguistic quality
As detailed in Section 6, we evaluated the linguistic quality of MAGIC contexts along two dimensions: naturalness and realism. The detailed evaluation criteria are provided in Figures 18 and 19.
@@
Appendix C Human Baseline Performance
To gain insight into how well humans can perform conflict detection and localization, we conducted a small-scale pilot study. Specifically, three trained crowd annotators independently reviewed 20 randomly sampled examples from MAGIC each.
As shown in Table 8, the annotators achieved an average of 92.5% for conflict identification (ID) and 83.3% for conflict localization (LOC), generally succeeding in identifying conflicts and their locations with reasonable accuracy.
While these results suggest that humans can reliably detect and localize conflicts–achieving substantially higher accuracy than current models (see Table 2 and Table 3)–they are based on a small-scale pilot and should not be taken as definitive. A more rigorous study with a larger and more diverse annotator pool will be conducted in future work.
Method ID (%) LOC (%)
Human Performance 92.5 83.3
Table 8: Human performance on the identification and localization evaluated on a small subset of MAGIC.